# Probing the Vulnerability of Large Language Models to Polysemantic Interventions

### v1

@@ H.2 Steering with Token Gradient Vector
H.3 Prompt Injection
In addition to the success rate of elevating target tokens to top-10 list, we evaluate the injection with other three metics. The first two ones are improvement in weighted cosine similarity and weighted overlap between model’s top-10 prediction and target token set. And the third one is improvement in best rank of target tokens in model’s top-10 prediction. We collect the best improvement for each sentence. The results of experiments on Pythia-70M, GPT-2-Small, LLaMA3.1-8B-Instruct and Gemma-2-9B-Instruct are shown below.13
Figure 13: Three metrics of Pythia-70M and GPT-2-Small on location and love&hate token set. GPT-2-Small uses green pillar and Pythia-70M uses blue pillar. The above two rows show three metrics evaluated on location type token set. The below two rows show three metrics evaluated on love&hate token set.
Figure 14: Three metrics of LLaMA3.1-8B-Instruct and Gemma-2-9B-Instruct on location and love&hate token set. LLaMA3.1-8B-Instruct uses purple pillar and Gemma-2-9B-Instruct uses orange pillar. The above two rows show three metrics evaluated on location type token set. The below two rows show three metrics evaluated on love&hate token set.
Appendix I Polysemantic Neuron Manipulation
@@
During the examination of SAE features and their connections with neurons, many features exhibit semantically similar activation texts. To avoid repetitive analysis on similar activation texts, we first perform feature clustering based on semantics of activation texts, and then check neuron connection at the cluster level. Given that the sparse auto-encoder from Neuronpedia is trained with a sparsity setting of $3$, the analysis focuses on the top three neurons with the highest alignment values per cluster. A threshold of $0.2$ is applied to filter out weak connections. Figure15 shows the distribution of polysemantic neurons identified in each layer. We can see that polysemantic neurons with strong connections with aggregated features only take up fewer than $5\%$ in each layer.
Figure 15: Distribution of polysemantic neurons in Pythia-70M.
Figure 16: Distribution of polysemantic neurons in GPT-2-Small.
@@
For strongly connected polysemantic neurons, we do further investigations on how suppressing or boosting their activation influence the semantic shift in model’s output to their aligned features. Neurons’ activation is multiplied with a scale value in the range $[0,20]$. Note that scaling within $[0,1]$ suppresses activation, while scaling within $[1,20]$ amplifies it. Results of more repetitive experiments are shown below17.
Figure 17: Replications of the neuron-level intervention on sampled neurons for GPT-2-Small.
Figure 18: Replications of the neuron-level intervention on sampled neurons for Pythia-70M.
Appendix J Ethics Statement, Limitations and Future Works
@@
This study has three key methodological limitations. First, we rely on SAEs to disentangle polysemantic activations; although SAEs are the de-facto tool, their outputs fluctuate with dimensionality and hyper-parameters, yielding unstable features (Paulo & Belrose, 2025; Heap et al., 2025; Gao et al., 2024). Second, our interventions steer only one interference feature in one layer, while multi-feature, cross-layer manipulations could amplify and better obscure the effect (Ameisen et al., 2025). Third, we quantify vulnerability solely via shifts in immediate next-token probabilities on two small base models—because only they both expose raw logits and have pre-trained SAEs—then check coarse transfer on two larger instructed models; establishing how these interventions alter non-trivial downstream tasks in bigger models is the next stage of this project.
@@
To balance reproducibility with responsible disclosure, we release complete code, evaluation scripts, and synthetic data, but deliberately omit the matrices that catalogue shared polysemantic directions. Publishing those mappings would make it easier to weaponize the very vulnerabilities we study, whereas the available artifacts still permit independent verification of all empirical claims.

### v2

@@ I.2 Experiment Details
K.2 Prompt Injection
Prompt injection in our experiments is done on the token types listed in Appendix E. Based on the classification of tokens, we also generate test sentences for each type, requiring that these sentences grammatically lead to tokens of the target type. The prompt is as follows.
System: You are a creative writing assistant specialized in generating incomplete sentences. Your task is to create incomplete sentences where the next logical token would very likely be from the category category.
Requirements:
1. Generate approximately 1000 incomplete sentences
2. Each sentence should be approximately 20 tokens long when complete
3. The sentences should end at a natural point where the next word would very likely be a category word
4. Use diverse contexts, scenarios, and grammatical structures
5. Make the sentences engaging and varied
6. Ensure the incomplete sentences create strong expectation for category words
Examples for category:
- For ’animal’: "In the dense jungle, we could hear the roar of a wild"
- For ’color’: "The sunset painted the sky a beautiful shade of"
- For ’emotion’: "When she heard the news, her face showed pure"
- For ’location’: "Our vacation destination this summer will be"
- For ’number’: "The recipe calls for exactly"
- For ’person’: "The award ceremony will be hosted by a famous"
- For ’science’: "The experiment required careful measurement of"
- For ’time’: "The meeting is scheduled for next"
Return your response as a Python list of strings, with exactly 1000 sentences. Format it properly as valid Python code that can be executed.
Start your response with: sentences = [
End your response with: ]
Do not include any explanatory text before or after the list.
User: Generate approximately 1000 incomplete sentences for the category category.
Available type tokens (separated by |, each token is enclosed in quotes):
type_token_1, type_token_2, ...
IMPORTANT NOTES:
- Each token above is a separate vocabulary item from language models
- Some tokens may have leading/trailing spaces (like " dog" or "cat ")
- These are the exact token strings that should be likely to appear as the next token after your incomplete sentences
- Consider the token boundaries when creating sentences
Please generate diverse, engaging incomplete sentences where the next word would very likely be from the category category tokens shown above. Make sure to use various contexts and grammatical structures.
Return as a Python list: sentences = [...]
The model returns approximately 400 to 800 sentences for each type. The dataset of token type denotation, along with the example sentences, has been made available in the GitHub repository.
After annotating the tokens in the model’s vocabulary with type labels, we partition the token sets based on the intervention information provided by the sparse auto-encoder. First, we filter for features where the highly activated tokens contain the target token type, simply tagging these as target features. Specifically, for each activation text, we identify the token with the highest activation value (max_act) and set a ratio (here, 0.8). Tokens with activation values exceeding ratio * max_act are considered highly activating the interference feature.
We then identify all interference features whose interference value with the target feature exceeds 0.2 and whose semantic similarity is below 0.3. Based on this set of interference features, we further partition them into high-interference and medium-interference sets. Features with interference values above a high_threshold (for a given target feature) are classified as high-interference features, while those with interference values between 0.2 and high_threshold are classified as medium-interference features.
Subsequently, we collect the highly-activating tokens from features in each set to form the respective token sets. All remaining tokens that were not collected constitute the random token set. It should be noted that the high-interference and medium-interference token sets exhibit significant overlap. To address this, we deduplicate the two sets to obtain disjoint token sets. For example, in the case of the Pythia-70M model, we annotate 1,938 tokens as belonging to the “location” type. Using the method described above and setting 0.5 as the high-interference threshold, the resulting high-interference token set contains 10,185 tokens, while the medium-interference set consists of 34,946 tokens. And 13,323 tokens remain in the random set. There is an overlap of 9,535 tokens between the high- and medium-interference sets. After deduplication, the high-interference set retains 650 tokens, and the medium-interference set retains 25,404 tokens.
For experiments on Pythia-70M and GPT-2-Small, we use the token sets generated by each respective model. For experiments on Llama-3.1-8B/70B-Instuct and Gemma-2-9B-Instruct models, we adopt the union of the token sets from the corresponding interference levels of the two small models. We examine a total of eight categories of token types: location, person, emotion, color, animal, science, number, and time. Apart from the three types mentioned in the main text that demonstrate strong generalizability, the test results for the remaining types are presented below.
Table 5: Token Types without Strong Generalizability
Target Model Original High-interference Low-interference Random
Person Pythia-70M $60.28\%***$ $29.02\%$ $28.58\%$ $32.31\%$
GPT-2-Small $54.29\%***$ $27.47\%***$ $26.97\%**$ $25.16\%$
Llama-3.1-8B-Instruct $38.85\%***$ $19.20\%***$ $19.08\%***$ $16.86\%$
Gemma-2-9B-Instruct $43.27\%***$ $25.36\%$ $26.10\%$ $25.04\%$
Llama-3.1-70B-Instruct $46.27\%***$ $21.90\%***$ $21.13\%**$ $19.61\%$
Animal Pythia-70M $96.96\%***$ $28.49\%$ $30.75\%$ $29.83\%$
GPT-2-Small $86.11\%***$ $21.64\%$ $18.87\%$ $25.03\%$
Llama-3.1-8B-Instruct $67.07\%***$ $42.76\%*$ $40.56\%$ $40.89\%$
Gemma-2-9B-Instruct $26.09\%***$ $41.02\%***$ $38.50\%$ $38.47\%$
Llama-3.1-70B-Instruct $49.32\%***$ $32.80\%*$ $30.79\%$ $31.43\%$
Emotion Pythia-70M $61.74\%***$ $26.13\%***$ $21.11\%$ $20.68\%$
GPT-2-Small $58.84\%***$ $36.91\%***$ $5.22\%* 33$$3.85\%$
Llama-3.1-8B-Instruct $60.16\%***$ $27.94\%$ $27.45\%$ $29.37\%$
Gemma-2-9B-Instruct $56.55\%***$ $13.85\%$ $13.63\%$ $13.50\%$
Llama-3.1-70B-Instruct $51.51\%***$ $44.90\%***$ $63\% 40.40$$40\%$
Color Pythia-70M $97.31\%***$ $17.89\%$ $21.49\%$ $21.80\%$
GPT-2-Small $85.47\%***$ $5.57\% 33$$3.13\%$ $36.93\%$
Llama-3.1-8B-Instruct $76.76\%***$ $22.01\%$ $20.67\%$ $20.97\%$
Gemma-2-9B-Instruct $22.57\%***$ $13.58\%$ $16.10\%$ $15.67\%$
Llama-3.1-70B-Instruct $76.85\%***$ $19.08\%$ $17.88\%$ $18.44\%$
Time Pythia-70M $69.92\%***$ $33\%*** 45.54$$54\%***$ $41.33\%$
GPT-2-Small $58.48\%***$ $31.55\%***$ $27.72\%***$ $21.42\%$
Llama-3.1-8B-Instruct $51.41\%***$ $25.58\%$ $26.29\%$ $25.95\%$
Gemma-2-9B-Instruct $55.15\%***$ $25.10\%*$ $25.94\%***$ $23.62\%$
Llama-3.1-70B-Instruct $78.28\%***$ $36.87\%$ $36.68\%$ $37.71\%$
Note: Cell values show the success rate of elevating target-type tokens into the top 30 predictions. Gray-shaded rows indicate black-box interventions. Testing uses a shared token set from the two small models. ***, **, and * denote t-test significance at $p<0.001$, $p<0.01$, and $p<0.05$, respectively, vs. random baseline. High- and low-interference tokens lie in $[0.5,1.0]$ and $[0.2,0.5]$. Details in Appendix K.2.
Appendix L Covert Intervention on Hellaswag Testset
To further examine the impact of the interference vectors obtained in the aforementioned sections on the overall performance of the model, we conduct a rapid validation using the experimental results targeting the “location” type of interference in Llama-3.1-8B-Instruct. Given that the intensity of the interference vectors applied in the previous experiments is aimed at maximizing the disruptive effect—which likely caused substantial impairment to the model, we first reduce the scale to $0.25$ times its original value. Subsequently, we randomly select $500$ test samples from the HellaSwag validation set and evaluate each interference vector on them. Using accuracy as the evaluation metric, the experimental results demonstrate that, out of a total of $174$ interference vectors, $149$ decreased the accuracy, while 25 increased it. The baseline accuracy is $77.2\%$, and the average reduction in accuracy is $2.77\%$. Some examples that keep model performance, i.e. reduce accuracy fewer than 11%, while still having substantial intervention effects are listed below.
Table 6: Covert Intervention on Hellaswag Dataset
Intervention Feature Top Predicted Tokens
Description of precautions related to safety protection “The documentary crew disappeared while filming in remote areas of” Raw Intervention the 0.35 the 0.24 Papua 0.027 Papua 0.04 Nepal 0.016 Africa 0.035
Numerical data that may require ordering or sorting “She always dreamed of owning a small cafe in” Raw Intervention the 0.34 the 0.36 a 0.27 a 0.28 her 0.21 Paris 0.13
Appendix M Transferable Polysemantic Structure Exploration
M.1 Brainstorming hidden associations
In this section, we analyze shared interference patterns in GPT-2-Small and Pythia-70M, which we previously demonstrated transfer to larger models under intervention. First, we use DeepSeek-V3 and GPT-5-mini to annotate every selected feature pair with the following prompt instruction.
System:
You are an expert in analyzing neural network feature semantics. Given two SAE (Sparse Autoencoder) features, their explanations, and their top activation texts, determine if they are semantically related.
For each feature, you will see:
- An explanation describing what the feature captures
- 5 text segments where the feature activates most strongly. High-activation tokens are marked with <<<token>>> or <<<multiple tokens>>>
Analyze the explanations and the marked tokens/activation patterns to determine:
1. Are these two features semantically related? Consider any form of semantic relationship - including direct overlaps (capturing similar concepts, linguistic patterns, or contextual meanings) as well as higher-order associations (e.g., semantic priming, thematic relatedness, or complementary roles).
2. If related, provide a concise description (max 50 tokens) of their relationship.
Return your analysis in this exact JSON format:
"isRelated": true/false, "description": "brief description" or null
Examples:
- Features activating on different tenses of verbs: "isRelated": true, "description": "Both capture verbal expressions, one for past tense, other for present tense"
- Features for numbers vs. animals: "isRelated": false, "description": null
- Features for positive vs. negative emotions: "isRelated": true, "description": "Both capture emotional expressions with opposite valence"
- Features for “doctor” vs. “hospital”: "isRelated": true, "description": "Conceptually linked via medical domain (profession vs. location)"
- Features for “boat” vs. “sand”: "isRelated": true, "description": "Loosely associated through beach/marine context (object vs. terrain)"
User: Feature A: feature_a
Explanation: feature_a’s explanation
Top activations: texts that activates feature_a the most
Feature B: feature_b Explanation: feature_b’s explanation
Top activations: texts that activates feature_b the most
To compare the behavior of the two automatic annotators (DeepSeek-V3 and GPT-5-mini), we compute the share of feature pairs labeled as “related” and the simple percent agreement in two subsets. In the Pythia-70M subset, DeepSeek-V3 labels $10.8\%$ of pairs as related and GPT-5-mini $28.7\%$, with $80.3\%$ percent agreement. In the GPT-2-Small subset, the corresponding proportions are $3.7\%$ (DeepSeek-V3) and $18.4\%$ (GPT-5-mini), with $83.8\%$ agreement. Aggregating both subsets, DeepSeek-V3 labels $9.5\%$ of pairs as related and GPT-5-mini $26.8\%$, and the overall percent agreement was $80.9\%$. We define percent agreement as the proportion of pairs for which both annotators assign the same label. Figure 15 A-B shows the agreement matrices for the two annotators. Interestingly, the overall consistency between the two annotators is low (Pythia-70M: Cohen’s $k=0.408$; GPT-2-Small: Cohen’s $k=0.221$). Compared with DeepSeek-V3, GPT-5-mini labels significantly more feature pairs as related, suggesting greater sensitivity to latent semantic associations. Nevertheless, more than $70\%$ of pairs are still judged completely unrelated in the annotators’ combined judgment (i.e., both say unrelated).
Figure 14: GPT-5-mini and DeepSeek-V3’s annotation reports for shared interfered feature pairs. DS is the abbreviation of DeepSeek-V3 and G5 is the abbreviation of GPT-5-mini.
In Table LABEL:tab:semrels, we further report eight qualitative examples where the two annotators flag latent associations and contrasting cases where they do not.
Table 7: Eight examples of annotated feature pairs.
Category Feature pair Interpretation
Frame Feature A: Phrases and sentences that highlight systemic issues related to incarceration and its effects on individuals and families. Feature B: Quantitative information related to statistics and predictions. Incarceration content often appears alongside quantitative/statistical references. (from GPT-5-mini)
Orthography Feature A: Occurrences of the word “acher.” Feature B: References to specific researchers or authors in studies related to minimum wage. Both capture researcher/author name substrings (e.g., -acher / Wascher). (from GPT-5-mini)
Axiology Feature A: The term “dear” in emotional contexts related to relationships and feelings of affection. Feature B: Concepts related to the notion of the sacred. Both capture emotionally significant concepts (affection vs. sacredness) with deep personal or spiritual value. (from DeepSeek-V3)
Homography Feature A: References to playing cards and card-related concepts. Feature B: References to the authors or studies related to economic analysis. Surface lexical overlap: “card” as playing card vs “Card” (author name). Same token, different senses. (from GPT-5-mini)
No-relation Feature A: References to “New Guinea.” Feature B: File formats and file compression terminology. None
No-relation Feature A: Instances of the abbreviation “ob” or related terms indicating observational data or annotations. Feature B: A specific term related to a well-known ride-sharing company. None
No-relation Feature A: Terms related to turbidity and its measurement. Feature B: References to specific music artists or groups. None
No-relation* Feature A: Occurrences of the name “Beethoven” and related variations. Feature B: Words related to expressions of frustration or annoyance. None
M.2 Feature-pair comparison test
Let $s_{\mathcal{M}}(f_{i},f_{j})$ denote the semantic similarity (in $\mathcal{M}$) between features $f_{i}$ and $f_{j}$, and let $I(f_{i},f_{j})$ denote their interference value.
From the set of shared interfering feature pairs in GPT-2-Small and Pythia-70M satisfying
$s_{\mathcal{M}}(f_{i},f_{j})<0.4\quad\text{and}\quad I(f_{i},f_{j})>0.4,$
we uniformly sample pairs for testing. For each sampled pair $(f_{1},f_{2})$, we then search for a matched comparison pair $(f_{3},f_{4})$ from the same layer such that
$\bigl|\,s_{\mathcal{M}}(f_{3},f_{4})-s_{\mathcal{M}}(f_{1},f_{2})\,\bigr|\leq\varepsilon\quad\text{and}\quad I(f_{3},f_{4})<0.1,$
i.e., semantic similarity is tightly matched while interference is substantially lower. For Pythia-70M, $\varepsilon=0.01$; for GPT-2-Small, due to limited matching, $\varepsilon=0.05$. Eventually, we collect $3,800$ feature-pair combination for this comparative analysis.
We then present the two pairs $(f_{1},f_{2})$ and $(f_{3},f_{4})$ (order randomized) to GPT-5-mini, asking the model to choose the more related pair using the prompt below.
System:
You are an expert in interpretability of LLM SAE features. You will receive two feature pairs (pair_one, pair_two). For each pair (f, g), use their explanations and Top-5 highest-activation texts (with high-activation tokens marked with <<<token>>> or <<<multiple tokens>>>) to decide which pair shows higher latent relatedness.
Definition:
"Latent relatedness" means the two features in a pair likely reflect the same underlying concept or are functionally coupled. Consider direct overlaps (concepts, patterns, meanings) and higher-order associations (thematic/complementary roles).
Decision criteria:
1) Explanations: semantic consistency, paraphrase/synonymy, or complementary roles within the same domain; penalize opposite/disjoint concepts.
2) Top-5 texts: overlap in topics/contexts; alignment of <<<markers>>> pointing to the same phrase, slot, entity, or role; penalize disjoint or contradictory evidence.
Uncertainty handling:
If evidence from (1)-(2) is insufficient or conflicting, return "uncertain".
Output:
Return ONLY a JSON object with exactly these keys:
"more_related_pair": "pair_one" or "pair_two" or "uncertain"
"reason": brief English rationale ($\leq$ 80 words), citing 1–2 decisive cues (e.g., explanation alignment, shared topics, token markers).
Must be valid JSON (double quotes). No extra text.
User: PAIR ONE
Feature 1
Explanation: explanation of feature one
Texts: five activation texts of feature one
Feature 2
Explanation: explanation of feature two
Texts: five activation texts of feature two
PAIR TWO
Feature 1
Explanation: explanation of feature one
Texts: five activation texts of feature one
Feature 2
Explanation: explanation of feature two
Texts: five activation texts of feature two
Appendix N Polysemantic Neuron Manipulation
@@ 6. Ensure the incomplete sentences create strong expectation for category words
During the examination of SAE features and their connections with neurons, many features exhibit semantically similar activation texts. To avoid repetitive analysis on similar activation texts, we first perform feature clustering based on the semantics of activation texts, and then check neuron connection at the cluster level. Given that the sparse auto-encoder from Neuronpedia is trained with a sparsity setting of $3$, the analysis focuses on the top three neurons with the highest alignment values per cluster. A threshold of $0.2$ is applied to filter out weak connections. Figure 15 shows the distribution of polysemantic neurons identified in each layer. We can see that polysemantic neurons with strong connections with aggregated features only take up fewer than $5\%$ in each layer.
Figure 15: Distribution of polysemantic neurons in each model. A is the result of Pythia-70M, and B is the result of GPT-2-Small. Error bars represent $95\%$ confidence intervals.
@@
For strongly connected polysemantic neurons, we do further investigations on how suppressing or boosting their activation influences the semantic shift in the model’s output to their aligned features. Neurons’ activation is multiplied with a scale value in the range $[0,20]$. Note that scaling within $[0,1]$ suppresses activation, while scaling within $[1,20]$ amplifies it.
Appendix O Ethics Statement, Limitations and Future Works
@@
This paper follows the ICLR Code of Ethics. This study has three key methodological limitations. First, we rely on SAEs to disentangle polysemantic activations; although SAEs are the de-facto tool, their outputs fluctuate with dimensionality and hyper‑parameters, yielding unstable features (Paulo & Belrose, 2025; Heap et al., 2025; Gao et al., 2024). Second, our interventions steer only one interference feature in one layer, while multi-feature, cross-layer manipulations could amplify and better obscure the effect (Ameisen et al., 2025). Third, we quantify vulnerability solely via shifts in immediate next-token probabilities on two small base models—because only they both expose raw logits and have pre-trained SAEs—then check coarse transfer on three larger instructed models; establishing how these interventions alter non-trivial downstream tasks in bigger models is the next stage of this project.
Appendix P Reproducibility Statement
@@
To balance reproducibility with responsible disclosure, we release complete code, evaluation scripts, and synthetic data in this Github repository, but deliberately omit the matrices that catalogue shared polysemantic directions between two small models. Publishing those mappings would make it easier to weaponize the very vulnerabilities we study, whereas the available artifacts still permit independent verification of all empirical claims.